What is the heat released when aqueous N2H4 is added to 5.50*10^3 L of plant water that is 2.50*10^-4 M O2? Enter your answer in scientific notation.
1
Expert's answer
2017-12-13T11:57:07-0500
N2H4(aq) + 4OH-(g)  N2(g) + 4H2O(l) + 4e-
O2(g) + 2H2O(l) + 4e-  4OH-(aq)

N2H4(aq) + O2(g)  N2(g) + 2H2O(l) H = -622.2 kJ
1.375 mol x kJ
n(O2) = 5.50103 L  2.5010-4 mol/L = 1.375 mol
x = 1.375 mol  (-622.2 kJ) / 1 mol = -855.5 kJ